OpenStudy (anonymous):

what fractions are equivalent to -3/8

pooja195 (pooja195):

Any fraction that you multiply the same number to example: \[\huge~\rm~\frac{ -3 \times 2=-6 /2=-3}{ 8 \times 2=16 /2=8}\]

pooja195 (pooja195):

See what I mean?

OpenStudy (anonymous):

@pooja195 the choices are -6/16, 12/32, -8/3, and 9/-24

OpenStudy (anonymous):

wait would it be -8/3 as one of the choices?

pooja195 (pooja195):

Nope because we cannot flip the fraction to make it the same how many answers are we allowed to choose?

OpenStudy (anonymous):

oh all four ,but probably only two

pooja195 (pooja195):

I see seems like i already gave one away :P

pooja195 (pooja195):

A.-6/16 <---this one I gave away .-. B.12/32 C.-8/3 <---this we can eliminate because we know it's not right D. 9/-24 With what is left what do you think the answer is? And why?

OpenStudy (anonymous):

um 9/-24?

OpenStudy (anonymous):

Because when I divide it, it turns out to be -3/8, right?

pooja195 (pooja195):

Correct :) well done

pooja195 (pooja195):

And its not 12/32 Because -3*4=-12 8*4=32